Alan is 4 times Peter's age. 12 years ago, he was 7 times as old. What are their ages now?
Answer by fractalier(6550)

Call their ages NOW, a and p.
a = 4p
a-12=7(p-12)=7p-84
or
a = 7p - 72
Substituting in we get
4p = 7p - 72
-3p = -72
p = 24
a = 96